I need help with this question, please: what is the equation? The Problem: Maria's take home pay last week was $585. This was after 1/4 of her pay was taken out for insurance and taxes. How much was Maria's gross pay before the deductions for taxes and insurance?
Answer by Greenfinch(383)

Pay less 1/4 pay = take home pay = 585
so 3/4 pay = 585
so Pay = 4/3 x 585 = $780